Circular No. 6236

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995aa IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

on technical pan films obtained with the OCA 0.9-m Schmidt telescope of an apparent supernova located at α = 22^{h}59^{m}09^{s}.73, δ = −4^{o}30′57′′.3 (equinox 2000.0), which is 8′′.1 east and 1′′.7 south of the galaxy’s center (the galaxy itself of mag V ∼ 17.0). Pollas provides the following V magnitude estimates for SN 1995aa: Sept. 2.02 UT, 19.8; 16.93, 20.2; 17.96, 20.2. A faint condensation (mag ∼ 21.0) is visible at this location on the Palomar Sky Survey and Digital Sky Survey. A nearby foreground star (V ∼ 16.5) has position end figures 09^{s}.76, 31′59′′.2.

S. Benetti, European Southern Observatory (ESO), reports that inspection of a noisy CCD spectrogram (range 390.0–925.0 nm, resolution 3.5 nm), obtained on Sept. 25.05 UT with the ESO 2.2-m telescope (+ EFOSC2), confirms that SN 1995aa is a peculiar type-IIn supernova. Hα emission is marginally resolved (peak measured at 782.4 nm; FWHM = 4.5 nm). Narrow emission lines of [O III] at 495.9–500.7 nm and He I at 587.6 nm are also present in the supernova spectrum. If these lines are at rest in the frame of the galaxy, a redshift of 0.19 is deduced.

COMET C/1995 O1 (HALE-BOPP) H. Rauer, Observatoire de Paris-Meudon (OPM); D. Despois, Ob-

servatoire de Bordeaux; R. Moreno and G. Paubert, Institut de Radio Astronomie Millimétrique (IRAM), Granada; and N. Biver, D. Bockelée-Morvan, P. Colom, J. Crovisier, and L. Jorda, OPM, report observations of the CO J(2–1) line at 230 GHz with the IRAM 30-m telescope: “The line was already marginally present in the average of Aug. 16 and 23 data, with an area of 0.087 ± 0.018 K km s^{−1} (antenna temperature scale). For the average of Sept. 20 and 21 data, the line was blueshifted by 0.33 km/s and had an area of 0.041 ± 0.003 K km s^{−1} and a width of 0.40 km/s. This corresponds to a CO production rate of ∼ 2 × 10^{28} s^{−1}, assuming a CO rotational temperature of 10 K, as was measured in 29P/Schwassmann-Wachmann 1 (Crovisier et al. 1995, Icarus 115, 213). The CO J(1–0) line could not be detected at the 3σ level of 0.03 K km s^{−1}.”

Total visual magnitude estimates: Sept. 16.84 UT, 9.9 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 17.16, 10.1 (C. S. Morris, Pine Mtn. Club, CA, 0.26-m refl.); 21.17, 11.2 (Morris); 22.85, 11.1 (Carvajal, 0.44-m refl.); 23.15, 11.3 (Morris); 23.84, 9.8 (Carvajal); 24.19, 10.8 (Morris).

1995 September 25 Daniel W. E. Green